1.      Jose spends her income $40 on two goods. Good X and Good Y. The price of good X is $8 and price of good Y is $2.

(a)   Draw the budget constraint to show how Jose can spend his money on the two goods.

(b)   Calculate the slope of the budget line.

(c)    If the price of good X increase by $2 what will happen to the budget line?

(d)   If the price of good Y decreases by $1 what will happen to the budget line? 


1
Expert's answer
2022-01-19T10:19:00-0500

Solution:

a.). Budget constraint: M = PxX + PyY

40 = 8X + 2Y

The budget constraint is as below:

 


 

b.). Slope of the budget line:

The slope of the budget line is = "-\frac{Px}{Py} = -\frac{8}{2} = -4"


c.). New budget line:

40 = 10X + 2Y

With an increase in the price of good X, the consumer can purchase fewer quantities of X than before, given her budget. Therefore, with the rise in the price of good X, the budget line will shift to the left to reflect the new position on the budget line.


d.). New budget line:

40 = 8X + Y

With a fall in the price of good Y, the consumer will be able to purchase more of good Y with the given budget, and therefore, the budget line will shift above to a new position.